OpenStudy (anonymous):

Write the equation of a line in point slope form that has a slope of -3 over 4 and passes through the point (4, 5).

OpenStudy (nurali):

slope-intercept form; y=mx+b Lets start by using point-slope form; (Y-Y1)=m(X-X1) (x1,y1)=(4, 5) and slope=-3/4 Now plug in the information given; (y-5)=-3/4(x-4) 4y-20=-3x+12 4y=-3x+12+20 4y=-3x+32

OpenStudy (swag):

@Nurali Thank you for helping her....But @riahhroo I would like to note that you should post math questions in the math section :)

OpenStudy (anonymous):

thank you so much.

OpenStudy (nurali):

Anytime.